OpenStudy (anonymous):

find an equation of the tangent line to the graph of f at the given point. f(x)=√3x^2-2) point(3,5)

OpenStudy (anonymous):

you need the derivative as the first step

OpenStudy (anonymous):

ok. brb. gonna get it

OpenStudy (anonymous):

the derivative of \(\sqrt{g(x)}\) is \[\frac{g'(x)}{2\sqrt{g(x)}}\]

OpenStudy (anonymous):

in your case it is \[\frac{6x}{2\sqrt{3x^2-2}}\]

OpenStudy (anonymous):

replace \(x\) by 3 to find the slope then use the all mighty point slope formula

OpenStudy (anonymous):

i got (1/2)(3x^2-2)^-1/2 (6x)

OpenStudy (anonymous):

luckily you get a nice rational number as an answer,

OpenStudy (anonymous):

hold the phone don't be a slave to the power rule. you are right of course, but writing the derivative that way is useless if you want to evaluate it at a number you should recall how to take the derivative of the square root of something. while every one else taking the exam is rewriting in exponential notation etc etc it will be to your benefit to know \[\huge\frac{d}{dx}[\sqrt{g(x)}]=\frac{g'(x)}{2\sqrt{g(x)}}\]

OpenStudy (anonymous):

woah!

OpenStudy (anonymous):

very simple the derivative of the thing inside the radical goes on top, divide by twice the radical of the inside piece instantly you get \[f'(x)=\frac{6x}{2\sqrt{3x^2-2}}\]now you can find \[f'(3)\] easilyy

OpenStudy (anonymous):

with the exponential notation, you cannot evaluate the derivative

OpenStudy (anonymous):

the final answer is supposed to look like 9x-5y-2=0

OpenStudy (anonymous):

we now need \(f'(3)\) which is possible when written properly \[f'(3)=\frac{6\times 3}{2\sqrt{3\times 3^2-2}}\] \[=\frac{18}{2\sqrt{25}}=\frac{18}{10}=\frac{9}{5}\]

OpenStudy (anonymous):

now you are 75% done point slope formula gives \[y-5=\frac{9}{5}(x-3)\]and it is algebra from here on in

OpenStudy (anonymous):

don't forget two common and important derivatives, just memorize them \[\frac{d}{dx}\sqrt{x}=\frac{1}{2\sqrt{x}}\] and by the chain rule \[\frac{d}{dx}[\sqrt{g(x)}=\frac{g'(x)}{2\sqrt{g(x)}}\]